### v1

# Walk Wisely on Graph: Knowledge Graph Reasoning with Dual Agents via Efficient Guidance-Exploration

Abstract
Recent years, multi-hop reasoning has been widely studied for knowledge graph (KG) reasoning due to its efficacy and interpretability. However, previous multi-hop reasoning approaches are subject to two primary shortcomings. First, agents struggle to learn effective and robust policies at the early phase due to sparse rewards. Second, these approaches often falter on specific datasets like sparse knowledge graphs, where agents are required to traverse lengthy reasoning paths. To address these problems, we propose a multi-hop reasoning model with dual agents based on hierarchical reinforcement learning (HRL), which is named FULORA. FULORA tackles the above reasoning challenges by eFficient GUidance-ExpLORAtion between dual agents. The high-level agent walks on the simplified knowledge graph to provide stage-wise hints for the low-level agent walking on the original knowledge graph. In this framework, the low-level agent optimizes a value function that balances two objectives: (1) maximizing return, and (2) integrating efficient guidance from the high-level agent. Experiments conducted on three real-word knowledge graph datasets demonstrate that FULORA outperforms RL-based baselines, especially in the case of long-distance reasoning.
Introduction
Knowledge graphs (KGs) are designed to represent the world knowledge in a structured way. There are various downstream NLP tasks especially knowledge-driven services, such as query answering (Guu, Miller, and Liang 2015; Cui et al. 2019), relation extraction (Mintz et al. 2009; Reiplinger, Wiegand, and Klakow 2014) and dialogue generation (He et al. 2017). However, a significant proportion of KGs are severely incomplete, which constrains their efficacy in numerous tasks. Consequently, this study concentrates on automatic knowledge graph (KG) reasoning, also as known knowledge graph completion (KGC).
Over the past few years, several embedding-based models (Bordes et al. 2013; Lin et al. 2015) are proposed to focus on preserving the structural information in the KG and are effective for single-hop reasoning, but lack the necessary interpretability. To solve this problem, some works (Xiong, Hoang, and Wang 2017; Das et al. 2018) have considered introducing reinforcement learning framework to KG reasoning tasks, these approaches compose single-hop triplets into multi-hop reasoning chains. With the development of deep learning (DL) and reinforcement learning (RL), recent multi-hop works (Wang et al. 2019; Lv et al. 2020; Nikopensius et al. 2023; Xu et al. 2024) incorporate more advanced DL and RL techniques. For example, AttnPath (Wang et al. 2019) makes use of attention mechanism to force an agent to walk forward every step to avoid the agent stalling at the same entity node constantly.
Although existing multi-hop reasoning models have achieved impressive results, there is a noteworthy issue with these models that they only perform well when the agent’s reasoning path length is short. The drawback that the agent relies heavily on short reasoning chains is fatal in certain datasets, such as sparse KG. Figure 1 is an illustration of short direct path and long indirect path. It’s challenging for an agent to identify a short direct path just like (a) due to the sparsity which makes the relation $r_{2}$ do not exist. The optimal path for agent is $e_{1}\rightarrow e_{2}\rightarrow e_{4}\rightarrow e_{3}$ we called long indirect path instead of $e_{1}\rightarrow e_{2}\rightarrow e_{3}$. From this point, enhancing the agent’s long-distance reasoning ability is one solution to alleviate the multi-hop models’ poor performance in the sparse KG.
Figure 1: An illustrative example of (a) short direct path and (b) long indirect path. When no short direct path exists, the agent searches for a long indirect path.
However, the dimension of the discrete action space at each space is large (Das et al. 2018). As the path length increases, the choices the agent faced will grow exponentially. The most relevant to our work are CURL (Zhang et al. 2022) which proposed a dual-agent framework and mutual reinforcement rewards to assign one of the agents (GIANT) searching on cluster-level paths quickly and providing stage-wise hints for another agent (DWARF). Compared to the classic multi-hop model MINERVA (Das et al. 2018), CURL does improve long-distance reasoning ability. But CURL also has two drawbacks, which make CURL’s performance inconsistent. (1) Mutual reinforcement reward mechanism forces the low-level agent DWARF to adopt similar policies to the high-level agent GIANT, even if GIANT’s policies is not well enough. This may result in false-negative rewards to the intermediate actions which are reasonable. (2) At the early phase, DWARF and GIANT adopt a near-random policy which makes low training efficacy. For briefness, we call the high-level agent as GIANT and the low-level agent as DWARF like (Bai et al. 2022).
A number of efforts have been made to justify Transformer in the RL architecture.
In light of these challenges, we propose FULORA, a robust dual-agent framework for KG reasoning with taking full advantage of entity embedding and relation embedding in the KG. FULORA seamlessly makes the self-exploration and path-reliance trade-off of DWARF through a supervised learning method. This unique mechanism will enable DWARF to make its own decisions while receiving meaningful guidance from GIANT, instead of relying entirely on the command of GIANT. Moreover, with the aim to make better use of the entity embedding and relation embedding, FULORA introduces attention mechanism and dynamic path feedback to DWARF and GIANT respectively. Intuitively, GIANT searches a feasible path as soon as possible to guide DWARF to reduce the search space, while DWARF adopts diversified exploration policies in the constrictive search space to prevent over-dependence on GIANT. In this way, DWARF walking on the original KG has both excellent long distance reasoning ability and short direct path utilization ability.
Related Work
In line with the focus of our work, we provide a brief overview of the background and related work on knowledge graph embedding and multi-hop reasoning.
Knowledge Graph Embedding
Knowledge graph embedding (KGE) methods map entities to vectors in low-dimensional embedding space, and model relations as transformations between entity embeddings (Bai et al. 2022). Once we map them to low-dimensional dense vector space, we can use modeling methods to perform calculations and reasoning (Ma et al. 2018). Prominent examples include TransE (Bordes et al. 2013), TransR (Lin et al. 2015), ConvE (Dettmers et al. 2018) and TuckER (Balazevic, Allen, and Hospedales 2019). These models are equipped with a scoring function that maps any triplet $(e_{s},r_{q},e_{o})$ to a scalar score. For instance, TransE (Bordes et al. 2013) first encoded the entities and relations into latent vectors by following translational principle in point-wise Euclidean space. The translation assumption in TransE implies the equation that $||\mathbf{h}+\mathbf{r}-\mathbf{t}||\approx 0$. The embeddings of entities and relations are learnt by maximizing the score function so that the likelihood score is high for true triplets while low for false triplets.
Multi-hop Reasoning
The advancement of deep reinforcement learning (DRL) has led to an increased interest in utilizing DRL to address path-finding issues. The first groundbreaking work combining DRL and KG reasoning is DeepPath (Xiong, Hoang, and Wang 2017), which has an important inspiration for the subsequent derivative models. However, the method is constrained by the requirement to know the target entity in advance. MINERVA (Das et al. 2018), by contrast, does not need to know the target entity in progress but instead causes the agent to walk on the knowledge graph until it finds answer.
In recent years, many powerful neural networks have been developed. Some works consider the potential of utilizing these networks to generate walking policies. M-Walk (Shen et al. 2018) utilizes RNN to record the sequence trajectories obtained by the agent traversal and then optimize the discounted reward by a RL algorithm called Monte Carlo Tree Search (MCTS), to tackle the problem of sparse rewards. GRL (Wang et al. 2020) employs both generative adversarial net (GAN) and long short-term memory (LSTM) to generate new trajectory sequences if needed, which means the agent cannot only reason in the original graph, but sub-graphs can be automatically generated to extend the relational path until the target entities are found. SQUIRE (Bai et al. 2022) utilizes an encoder-decoder Transformer structure to map query and previous path sequence to a contextualized representation, and further use such representation to autoregressively decode the output path, token by token. These models have demonstrated satisfactory performance in short-distance reasoning scenarios. However, their performance in long-distance reasoning tasks has consistently been suboptimal.
The proposal of hierarchical reinforcement learning (HRL) (Li et al. 2018; Wen et al. 2020; Liu et al. 2023) has brought attention to the fact that mupltiple agents can walk on the graph simultaneously. CURL (Zhang et al. 2022) aligns with the RL formulation of MINERVA (Das et al. 2018), i.e., learning to walk and search for answer entities of a particular KG query in an end-to-end fashion. Nevertheless, previous multi-agent based KG reasoning models don’t strike a satisfying balance between guidance from other agents and self-exploration. Moreover, these models don’t utilize enough of entity embedding and relation embedding, which wastes some structural information of the KG. In contrast to the preceding methodologies, FULORA employs a distinctive training approach that seamlessly integrates supervised learning and reinforcement learning. This innovative approach effectively reconciles the inherent trade-off between self-exploration and path-guidance, a challenge commonly encountered by DWARF.
Preliminary
In this section, we commence with the problem definition of our work, followed by the introduction of environment representation.
Problem Definition
We formally define the research problem of this paper in the following part. The knowledge graph is defined as a directed graph $\mathcal{G}=\{\mathcal{E},\mathcal{R}\}$, where $\mathcal{E}$ is a set of all entities and $\mathcal{R}$ is a set of all relations. A link triplet $l$ consists of the source entity $e_{s}\in\mathcal{E}$, the target entity $e_{o}\in\mathcal{E}$ and the relation $r\in\mathcal{R}$, i.e., $l=(e_{s},r,e_{o})$. In the real world, a link triplet corresponds to a fact tuple. For instance, we can represent the fact whale is a mammal as $(whale,is\ a,mammal)$. We follow the definition in prior graph walking models, that is, query answering (Das et al. 2018; Zhang et al. 2022). In the applications such as searching and query answering, most problems are to infer another entity when we only know the source entity $e_{s}$ and the query relation $r_{q}$, which can be formed by an incomplete link triplet $l_{m}=(e_{s},r_{q},?)$, where ”?” indicates the target entity $e_{o}$ is unknown and needs to be found in the KG.
Environment Representation
As in the standard reinforcement learning paradigm, we divide the environment which is modeled as a Markov decision process (MDP) into four elements: state, action, transition and reward.
State. The state consists of two entities and query relation. Specifically, the two entities consist of a current entity $e_{t}$ and a source entity $e_{s}$. The current entity visited by the agent at time step $t$ can be considered as state-dependent information, it reflects the reasoning condition of the agent. The source entity and query relation are the global context shared by all states. Concretely, $s_{t}=(e_{t},e_{s},r_{q})\in\mathcal{S}$.
Action. In reinforcement learning, the agent interacts with the environment by taking action. Specific to the graph walking task, the agent makes an action $a_{t}\in\mathcal{A}$ by selecting an outgoing edge of graph $\mathcal{G}$. This process can be formed as $a_{t}=\{(r_{t+1},e_{t+1})|(e_{t},r_{t+1},e_{t+1})\in\mathcal{G}\}$. In order to grant the agent an option to terminate a search, the agent can choose to stay at each entity by adding a self-loop edge to every action.
Transition. Choosing a new action entails a transition of state. A transition function $\delta:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is defined by $\delta(s_{t},a_{t})=\delta((e_{t},e_{s},r_{q}),a_{t})$, which means that performing an action at a given state produces a new state. Most multi-hop reasoning picks a neighboring entity at random because it’s common for an entity to have multiple neighbors that are all connected to it via the same relation (Xiong, Hoang, and Wang 2017; Das et al. 2018).
Reward. In the default formulation, after the agent takes an action under a certain state if the corresponding entity is a correct target, a favorable reward $1$ is obtained, and conversely, an unfavorable reward 00 is obtained.
Methodology
In this section, we propose FULORA, an extended dual-agent framework for knowledge graph reasoning via efficient guidance-exploration. As illustrated in Figure 4, FULORA is able to improve the exploration efficiency of GIANT via dynamic feedback mechanism, so as to provide more reliable guidance for DWARF. In a corresponding manner, DWARF employs a supervised learning approach to achieve a balance between guidance and exploration. Subsequently, it aggregates messages emitted by all neighbors of the current entity via an attention mechanism. Next, we will delve into the specifics of each aforementioned components.
Entity Embeddings and Cluster Embeddings
Consistent with (Zhang et al. 2022), we employ TransE (Bordes et al. 2013) due to its efficiency in encoding the structural proximity information of KG to generate pre-trained entity embeddings, then we divide the original KG into $N$ clusters by utilizing K-means. In order for GIANT can easily walk on the cluster-level graph $\mathcal{G}^{c}$ while preserving the relation information of the original KG $\mathcal{G}$, we add a link to two clusters if there is at least one entity-level edge between them, named cluster mapping. An example of cluster mapping is shown in Figure 2.
Figure 2: An example of cluster mapping.
Policy Networks For GIANT and DWARF
In our model, we utilize a three-layer LSTM, enabling the agent to memorize and learn from the actions taken before. In contrast to previous models, the necessity arises to design the network separately in this case, given that two agents are walking on the KG. An agent based on LSTM encodes the recursive sequence as a continuous vector $\mathbf{h}_{t}$. Specifically, the hidden state embedding of GIANT is $\mathbf{h}^{c}_{t}$ while the hidden state embedding of DWARF is $\mathbf{h}^{e}_{t}$. Their initial hidden state is $\mathbf{0}$. In addition, we define an information sharing vector $\mathbf{I}_{t}=[\mathbf{h}^{e}_{t};\mathbf{h}^{c}_{t}]$ for GIANT and DWARF to share path information. To a certain extent, cluster-level paths are complementary to entity-level paths, as they ensure the sharing of essential path information from GIANT to DWARF.
For GIANT. We denote the current cluster embedding at time step $t$ by $\mathbf{c}_{t}\in\mathbb{R}^{2d}$. The action representation $\mathbf{a}_{t}^{c}$ is given by the cluster embedding itself, i.e., $\mathbf{a}_{t}^{c}=\mathbf{c}_{t}\in\mathbb{R}^{2d}$ because the action corresponds to the next outgoing cluster. The history embedding is updated according to LSTM dynamics:
$\mathbf{h}_{t}^{c}=\text{LSTM}_{c}(\mathbf{W}^{c}[\mathbf{h}_{t-1}^{c};\mathbf{I}_{t-1}],\mathbf{a}_{t-1}^{c}),$ (1)
where $\mathbf{W}^{c}\in\mathbb{R}^{2d\times 4d}$ is a projection matrix to maintain shape.
For DWARF. Different from GIANT walking on the cluster KG, DWARF encounters greater difficulties on the original KG. Firstly, an entity has several different aspects. For example, a university professor may be linked with professional relations like worksForUniversity or worksInState, and also family relations like spouse or father/mother. Secondly, It can be observed that cluster-level paths are typically shorter than entity-level paths. Consequently, DWARF must possess enhanced long-distance reasoning capabilities to identify an accurate target. Furthermore, for diverse query relations, it’s advisable for DWARF to prioritize relations and neighbors that are highly correlated with the query relations. So we introduce Graph Attention mechanism (Velickovic et al. 2017) into DWARF.
A single-layer feedforward neural network is employed to compute attention weight. Consider that the current entity is $e_{i}$, a collection of all neighbors is $\mathcal{N}(i)$. The attention weight from $e_{i}$ and its neighbor $e_{j}\in\mathcal{N}(i)$ is calculated as
$\alpha_{ij}=\text{LeakyReLu}(\mathbf{a}[\mathbf{W}\mathbf{e}_{i}^{\top};\mathbf{W}\mathbf{e}_{j}^{\top}]),$ (2)
where $\mathbf{W}\in\mathbb{R}^{d\times d}$ is a linear transformation matrix and $\mathbf{a}\in\mathbb{R}^{2d}$ is a weight vector shared across all entities. For $e_{i}$, we normalize the attention weights with SoftMax:
$\alpha_{ij}=\frac{e^{\alpha_{ij}}}{\sum_{k\in\mathcal{N}(i)}e^{\alpha_{ik}}}.$ (3)
Now we can compute the graph attention for the state:
$\mathbf{atn}_{i}=\sum_{k\in\mathcal{N}(i)}\alpha_{ik}\mathbf{W}\mathbf{e}_{k}.$ (4)
DWARF’s history embedding $\mathbf{h}_{t}^{e}$ can be obtained from
$\mathbf{h}_{t}^{e}=\text{LSTM}_{e}(\mathbf{W}^{e}[\mathbf{h}_{t-1}^{e};\mathbf{atn}_{t-1};\mathbf{I}_{t-1}]),\mathbf{a}_{t-1}^{e}),$ (5)
where $\mathbf{W}^{e}\in\mathbb{R}^{2d\times 5d}$ is a projection matrix, while the action representation $\mathbf{a}_{t}^{e}$ is the concatenation of the relation embedding $\mathbf{r}_{t}\in\mathbb{R}^{d}$ and the end node embedding $\mathbf{e}_{t}\in\mathbb{R}^{d}$, i.e., $\mathbf{a}_{t}^{e}=[\mathbf{r}_{t};\mathbf{e}_{t}]\in\mathbb{R}^{2d}$.
Policy Generation.  To predict the next cluster for GIANT and the next entity for DWARF, we apply a two-layer feedforward network on the concatenation of their last LSTM states and current RL state embeddings,
$\begin{array}[]{ll}&\mathbf{d}^{c}_{t}=\text{SoftMax}(\mathbf{A}_{t}^{c}\times\mathbf{W}_{2}^{c}\text{ReLU}(\mathbf{W}^{c}_{1}[\mathbf{c}_{t};\mathbf{h}_{t}^{c}])),\\ &a_{t}^{c}\sim\text{Categorical}(\textbf{d}^{c}_{t}),\end{array}$ (6)
$\begin{array}[]{ll}&\mathbf{d}^{e}_{t}=\text{SoftMax}(\mathbf{A}_{t}^{e}\times\mathbf{W}_{2}^{e}\text{ReLU}(\mathbf{W}^{e}_{1}[\mathbf{e}_{t};\mathbf{r}_{q};\mathbf{h}_{t}^{e}])),\\ &a_{t}^{e}\sim\text{Categorical}(\textbf{d}^{e}_{t}),\end{array}$ (7)
where $\mathbf{W}_{1}^{c},\mathbf{W}_{2}^{c}\in\mathbb{R}^{4d\times 4d}$ and $\mathbf{W}_{1}^{e},\mathbf{W}_{2}^{e}\in\mathbb{R}^{6d\times 6d}$ are the matrices of learnable weights to maintain dimension of history embedding. While $\mathbf{A}_{t}^{c}\in\mathbb{R}^{|A_{t}^{c}|\times 4d},\mathbf{A}_{t}^{e}\in\mathbb{R}^{|A_{t}^{e}|\times 6d}$ represent the embeddings of all next possible actions for GIANT and DWARF respectively.
Efficient Guidance-Exploration
As mentioned above, in order to alleviate the issue of sparse rewards and huge action space, we hope GIANT will guide DWARF to shrink the action space. Nevertheless, the guidance provided by GIANT is not always beneficial, primarily due to two key issues: (1) Poor guidance will force DWARF to get wrong answer, and (2) even if GIANT is capable of generating fair guidance, mapping from entity to cluster results in a distribution shift, that is, a appropriate policy for GIANT may not be fully appropriate for DWARF.
Figure 3: An illustration of Effective Guidance-Exploration. When DWARF is out of bounds, GIANT guides it to move quickly inside. Otherwise, DWARF prefers to explore for itself to find a correct target.
Considering the two issues, we propose an efficient guidance-exploration approach, which gives DWARF a constraint reward to balance guidance and exploration via supervised learning approaches.
Constraint Reward.  Considering the constraint, that is, receiving high-quality guidance from GIANT, we introduce a metric for the state distance between GIANT and DWARF, denoted as $D(s_{t}^{c},s_{t}^{e})$. It is calculated as the cosine similarity between the pre-trained embeddings of the current entity and the current cluster:
$D(s_{t}^{c},s_{t}^{e})=\frac{\mathbf{c}_{t}^{\top}\mathbf{e}_{t}}{||\mathbf{c}_{t}||_{2}||\mathbf{e}_{t}||_{2}}.$ (8)
We solve the optimization problem by the following formulas,
maximize $\text{maximize }\mathbb{E}_{a_{1}^{e},...,a_{T}^{e}\sim\pi_{\theta}^{e}}\left[\sum_{t=0}^{T-1}r_{e}(s_{t}^{e}|s_{0}^{e})\right]$ (9)
subject to $\text{subject to }D(s_{t}^{c},s_{t}^{e})\geq\frac{\delta}{r_{c}(s_{t}^{c})+\varepsilon}$
Figure 4: An overview of FULORA framework.
where $r_{e}(s_{t}^{e})$ and $r_{c}(s_{t}^{c})$ are default rewards for DWARF and GIANT respectively, the agent obtains a favorable reward 1 if the corresponding entity or cluster is a correct target and unfavorable reward 0 otherwise. It is only when GIANT reaches the correct cluster that the constraint in Equation 9 is operative.
Practical Algorithm.  The objective in Equation 9 can be optimized with any reinforcement learning algorithm that implements generalized policy iteration. Here we use REINFORCE (Williams 1992) and the method of Lagrange multipliers as described by (Abdolmaleki et al. 2018; Grillotti et al. 2024). For all DWARF state $s_{t}^{e}$, we maximize the Lagrangian function, subject to $0\leq\lambda(s_{t}^{e})\leq 1$,
$J(\theta_{\pi^{e}})=\sum_{t=0}^{T-1}[(1-\lambda(s_{t}^{e}))r_{e}(s_{t}^{e})+\lambda(s_{t}^{e})D(s_{t}^{c},s_{t}^{e})],$ (10)
the Lagrange multiplier is updated to make the guidance-exploration trade-off. When the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is less than the threshold, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ increases to encourage GIANT to provide more guidance for DWARF. Conversely, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ decreases to encourage DWARF to explore in the constrictive space when the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is greater than the threshold. In practice, we utilize a cross-entropy loss to optimize $\theta_{\lambda}$:
$J(\theta_{\lambda})=\sum_{t=0}^{T-1}[-(1-y)\text{log}(1-\lambda(s_{t}^{e}))-y\text{log}(\lambda(s_{t}^{e}))]$ (11)
$\text{where }y=\left\{\begin{aligned} &0\quad\quad\text{if}\ D(s_{t}^{c},s_{t}^{e})\geq\frac{\delta}{r_{c}(s_{t}^{c})+\varepsilon}\\ &1\quad\quad\text{otherwise}\end{aligned}\right.$
Dynamic Path Feedback
A further challenge is to enhance the search efficiency of GIANT in order to provide DWARF with the requisite guidance as expeditiously as possible, given that the search is limited to a fixed number of step $T$. In the default reward, GIANT will receive 1 reward only when it reaches the correct cluster, and the rest will be 0, which makes GIANT adopt random policy at the early phase. This phenomenon is not conducive to stable learning outcomes for two agents. On the basis of the theory of reward shaping (Ng, Harada, and Russell 1999; Harutyunyan et al. 2015), we rewrite the reward function of GIANT, named dynamic path feedback. In particular, we write it in the form of an objective function $J(\theta_{\pi^{c}})$,
$J(\theta_{\pi^{c}})=\sum_{t=0}^{T-1}[r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})],$ (12)
$\text{where}\ \Delta(s_{t}^{c},s_{t+1}^{c})=D(s_{t}^{c},s_{\text{target}}^{c})-D(s_{t+1}^{c},s_{\text{target}}^{c})$
$s_{t+1}^{c}$ comes from the next state generated by the policy network. In contrast to the default reward, dynamic path feedback uses the reward function to score the GIANT’s path in a rollout rather than simply identifying whether it has reached the correct target cluster. Even if GIANT does not reach the correct target cluster in a rollout, it will evaluate the quality of the path, thereby accelerating the learning process. In Appendix D, we prove that GIANT learns optimal policy in dynamic path feedback is consistent with default rewards.
Algorithm 1 FULORA Training Algorithm
0:    KG $\mathcal{G}$; Initial policy networks parameters $\theta_{\pi^{e}}$ and $\theta_{\pi^{c}}$; Initial Lagrange multiplier parameter $\theta_{\lambda}$ ;Source entity and cluster nodes $e_{s}$ and $c_{s}$; Entity-level query $r_{q}$; Target entity and cluster nodes $e_{o}$ and $c_{o}$; Maximum path length $T$
0:    parameters $\theta_{\pi^{e}}$, $\theta_{\pi^{c}}$, $\theta_{\lambda}$
1:  for $t=0,...,T-1$ do
2:     Set default cluster-level reward $r_{c}=1$ if $c_{t}=c_{o}$ otherwise $r_{c}=0$
3:     Set default entity-level reward $r_{e}=1$ if $e_{t}=e_{o}$ otherwise $r_{e}=0$
4:     Predict the action $a_{t}^{c}$ and $a_{t}^{e}$ for GIANT and DWARF based on policy networks parameters $\theta_{\pi^{c}}$ and $\theta_{\pi^{e}}$
5:     Compute $J(\theta_{\pi^{c}})$, $J(\theta_{\pi^{e}})$, $J(\theta_{\lambda})$ based on Eq. (10)-(12).
6:  end for
7:  Update model parameters: $\theta_{\pi^{c}}\leftarrow\theta_{\pi^{c}}+\alpha_{\pi^{c}}\nabla_{\theta_{\pi^{c}}}J(\theta_{\pi^{c}})$ $\theta_{\pi^{e}}\leftarrow\theta_{\pi^{e}}+\alpha_{\pi^{e}}\nabla_{\theta_{\pi^{e}}}J(\theta_{\pi^{e}})$ $\theta_{\lambda}\leftarrow\theta_{\lambda}-\alpha_{\lambda}\nabla_{\theta_{\lambda}}J(\theta_{\lambda})$
8:  return  $\theta_{\pi^{e}}$, $\theta_{\pi^{c}}$, $\theta_{\lambda}$
Experiments
In this section, we evaluate the efficacy of FULORA on three real-world KG datasets: NELL-995 (Xiong, Hoang, and Wang 2017), WN18RR (Dettmers et al. 2018) and FB15K-237 (Toutanova et al. 2015). The datasets statistics are listed in Table 1.
Dataset #Ent #Rel #Fact #Que #$\textbf{\#Deg}_{\textbf{mean}}#$$\textbf{\#Deg}_{\textbf{med}}#$
NELL-995 75,492 200 154,213 3,992 4.07 1
WN18RR 40,945 11 86,835 3,134 2.19 2
FB15K-237 14,505 237 272,115 20,466 19.74 14
Table 1: The statistics of some benchmark KG datasets. #Deg is the outgoing degree of every entity that can indicate the sparsity level.
As can be seen from the statistical indicators, these three datasets represent standard KG, sparse KG and Dense KG respectively. In our selection of the baseline, we are not only evaluating it against the state-of-the-art multi-hop reasoning methods (DeepPath (Xiong, Hoang, and Wang 2017), MINERVA (Das et al. 2018), M-Walk (Shen et al. 2018), AttnPath (Wang et al. 2019), SQUIRE (Bai et al. 2022) and CURL (Zhang et al. 2022)), but also against with other embedding-based KG reasoning methods (TransE (Bordes et al. 2013), DistMult (Yang et al. 2015), ComplEx (Trouillon et al. 2016) and LMKE (Wang et al. 2022)). FULORA and all baselines are implemented under the Pytorch framework and run on the NVIDIA 3080Ti GPU. We set the default path length to 3, and all the results are the average of the results in five experiments.
Link Prediction Results
For each triplet $(e_{s},r_{q},e_{o})$ in the test set, we convert it to a triplet query $(e_{s},r_{q},?)$, and then use embedding-based models or multi-hop based models and a beam search width of 50 to get the ranking list of tail entity. Following the previous work (Bordes et al. 2013), here we use two metrics: (1) the mean reciprocal rank of all correct tail entities (MRR), and (2) the proportion of correct tail entities ranking in the top K (Hits@K) for evaluation.
Model NELL-995 WN18RR FB15K-237
MRR @1 @3 @10 MRR @1 @3 @10 MRR @1 @3 @10
TransE 51.4 45.6 67.8 75.1 35.9 28.9 46.4 53.4 36.1 24.8 40.1 45.0
DistMult 68.0 61.0 73.3 79.5 43.3 41.0 44.1 47.5 37.0 27.5 41.7 56.8
ComplEx 68.4 61.2 76.1 82.1 41.5 38.2 43.3 48.0 39.4 30.3 43.4 57.2
LMKE 74.6 71.7 84.7 89.5 53.6 45.1 66.0 79.4 41.2 31.8 46.2 56.9
MINERVA 67.5 58.8 74.6 81.3 44.8 41.3 45.6 51.3 27.1 19.2 30.7 42.6
M-Walk 70.7 63.2 75.7 81.9 43.7 41.5 44.7 54.3 23.4 16.8 24.5 40.3
AttnPath 69.3 62.7 73.9 80.1 42.9 40.7 44.3 52.9 31.9 24.1 40.4 43.8
SQUIRE 71.1 68.2 81.4 87.2 48.2 45.0 51.4 59.7 37.4 28.0 42.1 52.5
CURL 70.8 66.7 78.6 84.3 46.0 42.9 47.1 52.3 30.6 23.9 38.1 50.9
FULORA 72.5 69.4 79.7 89.2 49.1 45.6 50.7 57.2 33.4 25.4 39.2 50.7
Table 2: Link prediction results on three datasets from NELL, WordNet and Freebase. All metrics are multiplied by 100. The best score of multi-hop reasoning models is in bold, and the best score of embedding-based models is underlined.
As shown in Table 2, we first display the performances of FULORA and all baselines on NELL-995, WN18RR and FB15K-237. From the table, we can observe that our model outperforms all previous multi-hop reasoning models on standard KG (NELL-995) and sparse KG (WN18RR). Especially on NELL-995, where our model gains marked improvements compared with the best multi-hop reasoning baseline SQUIRE and dual-agent model CURL. In addition, FULORA scores similarly to the best embedding-based model LMKE. Since FB15K-237 contains much larger number of 1-to-M than the M-to-1 relation instances (Wan et al. 2020; Zhang et al. 2022), multi-hop reasoning methods is prone to be stuck in the local entity nodes with high-degree centrality, renders it hard to reach the correct entity. Compared with strigent multi-hop reasoning models (MINERVA, M-Walk and CURL), FULORA’s attention mechanism alleviates this issue by forcing the agent to focus on the neighbor which is highly relative to the query. After averaging results on three datasets, we observe that FULORA leads to overall improvements relative to the dual-agent model (CURL) by 2.5%, 2.3%, 1.3%, 3.2%, in terms of MRR and Hits@1, 3, 10.
Fact Prediction Results
As opposed to link prediction, fact prediction task is concerned with verifying the veracity of an unknown fact, the true test triplets are ranked with some generated false triplets. Since we share a similar query-answering mechanism as CURL (Zhang et al. 2022), FULORA is capable of identifying the most appropriate entity for a given query and eliminates the need to evaluate negative samples of any particular relation. In the experiments, the dual agents try to infer and walk through the cluster-level and entity-level respectively to reach the correct target under removing all links of groundtruth relations in the original KG. Here, we report Mean Average Precision (MAP) scores for various relation tasks of NELL-995. We reuse the results of TransE, TransR, PRA, DeepPath, MINERVA, M-Walk on seven tasks already reported in (Zhang et al. 2022). As demonstrated in Table 3, FULORA produces a satisfying result in most tasks, contributing an average gain of 9.1% relative to the multi-hop based reasoning approaches (PRA, DeepPath, MINERVA, M-Walk and CURL) and 16.1% gain compared to the embedding-based approaches (TransE and TransR).
Long-Distance Reasoning Ability
Recall the motivation of FULORA, we tend to address the issue of multi-hop reasoning models being unable to infer correct answer due to the lack of short direct path by improving long-distance reasoning ability. This issue is significant on standard KG and sparse KG, so we conduct the following experiments on NELL-995 and WN18RR. For effective evaluation, we compare our model with MINERVA and CURL in NELL-995 and WN18RR, where we remove the most frequently-visited short paths found by the bi-directional search (Xiong, Hoang, and Wang 2017) inside KGs.
Task TransE TransR PRA DeepPath MINERVA M-Walk CURL FULORA
PersonBornInLocation 62.7 67.3 54.7 75.1 80.0 84.7 82.7 84.4
OrgHeadquarteredInCity 62.0 65.7 81.1 79.0 94.0 94.3 94.8 95.1
AthletePlaysForTeam 62.7 67.3 54.7 75.0 80.3 84.7 82.9 86.0
AthletePlaysInLeague 77.3 91.2 84.1 96.0 94.2 96.1 97.1 97.3
AthletePlaysSport 87.6 96.3 47.4 95.7 98.0 98.3 98.4 98.5
TeamPlaysSport 76.1 81.4 79.1 73.8 88.0 88.4 88.7 90.2
WorksFor 67.7 69.2 68.1 71.1 81.0 83.2 82.1 84.3
Table 3: Fact prediction results on seven tasks from NELL-995. All metrics are multiplied by 100. The best score of all models is in bold.
Figure 5: Learning curves comparing the performance of ours against CURL from NELL-995 relation tasks and all tasks. Averaged over 5 seeds with the shaded area showing standard deviation. Our proposed model is significantly better than CURL in both score and stability.
Cluster State Similarity.   One of the most significant contributions of GIANT is the implementation of dynamic path feedback, which has been shown to enhance the learning efficiency. To visually demonstrate the efficacy of dynamic path feedback, we initially focus on GIANT’s reasoning ability on cluster-level, which directly impacts DWARF’s reasoning results. Here, we record the cluster state similarity (CSS) $D(s_{T}^{c},s_{\text{target}}^{c})$ under each epoch. As shown in Figure 5, our model outperforms CURL in scores and stability in most cases. This demonstrates that dynamic path feedback does improve GIANT’s learning efficiency. See Appendix B for a comparison of the remaining tasks.
Entity Reasoning Accuracy.  The preceding analysis has demonstrated that FULORA enables GIANT to provide high-quality guidance, and next, we will demonstrate that efficient guidance-exploration gives DWARF a superior performance in long-distance reasoning against other multi-hop baselines. Figure 6 plots the MRR scores on varying path lengths on NELL-995 and WN18RR. As can be seen, FULORA outperforms CURL, MINERVA and SQUIRE with different path lengths. In particular, FULORA is much more prominent in the long-distance reasoning accuracy of WN18RR (sparse KG). It explains our motivation: FULORA can alleviate the poor performance of multi-hop reasoning approaches in the absence of a short direct path (typically found in standard KG and sparse KG) by improving long-distance reasoning ability. Overall, FULORA has a more robust performance under the demands of long-distance reasoning, and also shows minimal performance degradation for long-path settings. It is mainly due to the dynamic path feedback and efficient guidance-exploration proposed by us, which enable more efficient information sharing between GIANT and DWARF, and ensure that DWARF maintains an excellent self-exploration ability.
Figure 6: The long-distance performance: FULORA significantly outpeforms CURL, MINERVA, SQUIRE on NELL-995 and WN18RR.
Ablation Study
We utilize $\delta$ and $\alpha$ to control the degree of efficient guidance-exploration and dynamic path feedback. To comprehensively explore the efficacy of efficient guidance-exploration and dynamic path feedback, we conduct link prediction on NELL-995 and WN18RR by varying the values of $\delta$ and $\alpha$ as $\{0.20,0.30,0.40,0.50\}$ and $\{0.05,0.10,0.15,0.20\}$, respectively. The averaged results are shown in Figure 7. In accordance with our previous analysis, it is evident that DWARF cannot rely excessively on GIANT for guidance. Furthermore, the level of path feedback that GIANT receives should not be unduly strong. In addition, in Appendix B, we have carefully discussed the influence of three main components of FULORA on knowledge graph reasoning efficiency.
Figure 7: Ablation analysis w.r.t $\delta$ and $\alpha$ on NELL-995 and WN18RR, respectively. Here we report MRR results with path length=3 for both datasets.
Conclusion
We present FULORA, an efficient guidance-exploration model built on dual-agent KG reasoning framework to enhance the agent’s long-distance reasoning ability on standard KG and sparse KG. The key insight behind our approach is balancing the self-exploration of DWARF and the guidance from GIANT. Specifically, on the one hand, we leverage the attention mechanism to make DWARF pay attention to the neighbouring entities that are close to the query. On the other hand, we propose that dynamic path feedback enables GIANT to have better learning efficiency, thus providing DWARF with high-quality guidance, making the DWARF to have a favourable global vision while having excellent local reasoning ability. Experiments on three real-world datasets show that FULORA outperforms the recent state-of-the-art multi-hop reasoning methods in most cases. Further analysis of these results reveals that FULORA’s Long-distance reasoning ability on standard KG and sparse KG has significant advantages compared to multi-hop reasoning methods.
Appendix A A. Experiment Setup
Data Statistics
We adopt three knowledge graph datasets with different scales: NELL-995, WN18RR and FB15K-237. We give a brief overview of these datasets.
NELL-995 (Xiong, Hoang, and Wang 2017). NELL995 is an open source machine learning dataset developed by the OpenAI research group that contains more than 950,000 pieces of entity relationship data collected from the network to help machine learning systems make inferences. NELL995 can be used to train machine learning models, such as natural language processing models, machine translation models, question answering systems, and semantic search systems.
WN18RR (Dettmers et al. 2018). WN18RR is a subset of WordNet that describes the association characteristics between English words. It preserves the symmetry, asymmetry, and composition relationships of the WordNet, and removes the inversion relationships. WN18RR contains some relational information about words. It consists of 40,945 entities and 11 relations.
FB15K-237 (Toutanova et al. 2015). FB15K-237 is a common knowledge graph dataset, which is a subset extracted from Freebase knowledge graph. It contains 14,505 entities and 237 relations, and the data is carefully processed to remove reversible relational data and trivial triples, ensuring that entities in the training set are not directly connected to the verification or test set, thus avoiding information leakage issues.
Baseline Methods
We compare the performance of FULORA with the following baselines, including embedding-based and multi-hop-based knowledge graph reasoning methods:
  1. 1.
    TransE (Bordes et al. 2013) embeds entities and relations into low-dimensional vector space so that complex relationships in graphs can be represented and reasoned by vector operations.
  2. 2.
    DistMult (Yang et al. 2015) uses triples (source entity, relation, tail entity) to train the model, where each entity and relation is represented as a vector. The target is to maximize the score of the correct triples while minimizing the score of the wrong triples.
  3. 3.
    ComplEx (Trouillon et al. 2016) is based on tensor decomposition and uses complex vectors to represent entities and relationships in order to capture complex relationships in the knowledge graph.
  4. 4.
    DeepPath (Xiong, Hoang, and Wang 2017) utilizes a knowledge graph-based embedded policy-based agent with continuous states, which extends its path in a KG vector space by sampling the most promising relationships.
  5. 5.
    MINERVA (Das et al. 2018) formulates the query task as a reinforcement learning (RL) problem where the goal is to take the best sequence of decisions (choice of relation edges) to maximize the expected reward (reaching the correct answer node).
  6. 6.
    M-Walk (Shen et al. 2018) uses RNNs to convert the historical path into a vector for learning policy and Q-function. At the same time, because of the sparsity of reward, RNN with Monte Carlo tree search is used to generate paths.
  7. 7.
    AttnPath (Wang et al. 2019) makes use of attention mechanism to force an agent to walk forward every step to avoid the agent stalling at the same entity node constantly.
  8. 8.
    SQUIRE (Bai et al. 2022) utilizes an encoder-decoder Transformer structure to translate the query to a path, without relying on existing edges to generate the path.
  9. 9.
    LMKE (Wang et al. 2022) adopts Language Models to derive Knowledge Embeddings (KE) and formulate description-based KE learning with a contrastive learning framework to improve efficiency in training and evaluation.
  10. 10.
    CURL (Zhang et al. 2022) trains two agents (GIANT and DWARF) to walk over a knowlegde graph jointly and search for the answer collaboratively.
Implementation Details
We use the following software versions:
  • Ubuntu 24.04 LTS
  • Python 3.10
  • Pytorch 2.0.1
We conduct all experiments with a single NVIDIA GeForce 3080Ti GPU.
Our experiments of all of 10 baselines are conducted on their official implementation provided by their respective authors. To reproduce the results of our model in Table 2 and Table 3, we report the empirically optimal crucial hyperparameters as shown in Table 4. On all datasets, the quantities of path rollouts in training and testing are 20 and 100, separatively.
Attribute NELL-995 WN18RR FB15K-237
Embedding size 50 50 50
Hidden size 50 50 50
Batch size 128 256 512
Learning rate 0.01 0.001 0.001
Optimizer Adam Adam Adam
Cluster number 75 100 200
Beam search size 100 50 100
$\alpha$ 0.15 0.15 0.25
$\delta$ 0.20 0.40 0.30
Table 4: FULORA Hyperparameters on three KG datasets.
Appendix B B. Additional Results and Analysis
Case Stuides
WorksFor (Answer):  Jeff Skilling $\overset{worksfor}{\rightarrow}$ Enron
WorksFor (AttnPath):  Jeff Skilling(2) $\overset{worksfor}{\rightarrow}$ Enron(2) $\overset{topmemberoforganizition}{\rightarrow}$ Kenneth Lay $\overset{personleadsorganization}{\rightarrow}$
Enron and Worldcom(5)
WorksFor (CURL):  Jeff Skilling(2) $\overset{personleadsorganization}{\rightarrow}$ Enron and Worldcom
$\overset{subpartoforganization}{\rightarrow}$ GE $\overset{subpartoforganization^{-1}}{\rightarrow}$ Enron and Worldcom $\overset{personleadsorganization^{-1}}{\rightarrow}$ Jeff Skilling
$\overset{worksfor}{\rightarrow}$ Enron(3)
WorksFor (FULORA):  Jeff Skilling(4) $\overset{worksfor}{\rightarrow}$ Enron $\overset{topmemberoforgnazition}{\rightarrow}$ Kenneth Lay(2)
$\overset{worksfor}{\rightarrow}$ Enron(3)
AthletePlaysSport (Answer):  Carlos Villanueva $\overset{athleteplayssport}{\rightarrow}$ Baseball
AthletePlaysSport (AttnPath):  Carlos Villanueva(2) $\overset{athleteflyouttosportsteamposition}{\rightarrow}$ Center
$\overset{athleteflyouttosportsteamposition^{-1}}{\rightarrow}$ Chris Coste(3) $\overset{athleteplayssport}{\rightarrow}$ Baseball(2)
AthletePlaysSport (CURL):  Carlos Villanueva(4) $\overset{athleteflyouttosportsteamposition}{\rightarrow}$ Center
$\overset{athleteflyouttosportsteamposition^{-1}}{\rightarrow}$ J.C. Boscan $\overset{athleteplayssport}{\rightarrow}$ Baseball(2)
AthletePlaysSport (FULORA):  Carlos Villanueva(3) $\overset{athleteplayssport}{\rightarrow}$ Baseball(5)
TeamPlaysinLeague (Answer):  Duke University $\overset{teamplaysinleague}{\rightarrow}$ International
TeamPlaysinLeague (AttnPath):  Duke University $\overset{templayssport}{\rightarrow}$ Basketball $\overset{teamplayssport^{-1}}{\rightarrow}$ Boise State Broncos
$\overset{teamplaysinleague}{\rightarrow}$ NCAA
TeamPlaysinLeague (CURL):  Duke University $\overset{teamplayssport}{\rightarrow}$ Basketball $\overset{teamplayssport^{-1}}{\rightarrow}$ Bucks
$\overset{teamplaysinleague}{\rightarrow}$ NBA
TeamPlaysinLeague (FULORA):  Duke University $\overset{teamplayssport}{\rightarrow}$ Basketball $\overset{teamplayssport^{-1}}{\rightarrow}$ Old Dominion University
$\overset{teamplaysinleague}{\rightarrow}$ International
Table 5: Paths found by AttnPath, CURL and FULORA, respectively. In order to describe the self-ring succinctly, we denote that agent stay at the same entity continuously $n$ times as Entity($n$).
[…]
Figure 8: Learning curves comparing the performance of ours against CURL. Learning curves are averaged over 5 seeds, and the shaded area represents the standard deviation across seeds.
Cluster State Similarity
To assess the exploration efficiency of GIANT, we introduce cluster state similarity (CSS) as a metric, and in the main content we compare the performance of CURL and FULORA on only three tasks due to space constraints. Figure 8 also compare the model performance on remaining six tasks. For the same task, CURL experiences significant oscillations as path length increases, while FULORA remains stable. As in our previous analysis, CURL can only judge path quality by whether it has gone to the correct target cluster, while FULORA utilizes dynamic path feedback to promote the GIANT to converge to a high-quality path.
Figure 9: The difference value to ratios of the mean and variance of CSS between FULORA and CURL on different task-path_length.
Figure 9 further visualizes performances of FULORA and CURL on these tasks with varied path length. We utilize the ratio of the mean to the variance of CSS as an evaluation metric for the mean measures inference accuracy and the variance measures inference stability. In the vast majority of cases, FULORA outperforms CURL. In the case of long-distance reasoning for complex tasks like PersonBorninLocation-Path9 (PBL-9), AthletePlaysinLeague-Path7 (APL-9), AthletePlaysSport-Path9 (APS-9), and TeamPlaysinLeague (TPL), the performance gap between FULORA and CURL is significant.
Additional Ablation Studies
Here, we conduct a series of ablation experiments designed to answer the following three questions:
Q1: Can attention mechanism enhance DWARF’s reasoning ability?
Q2: Can dynamic path feedback accelerate DWARF’S learning speed via improving GIANT’S reasoning ability?
Q3: Can the efficient guidance-exploration approach make the guidance-exploration trade-off?
With a slight abuse of abbreviation, we use ATTN, DPF and GE to denote attention mechanism, dynamic path feedback and the efficient guide-exploration approach respectively. To highlight the effect of each component, we set FULORA as the baseline. In Figure 10, FULORA-X represents pulling module X out of the FULORA framework. We conduct experiments on three different scales of KG to make a fair comparison. It is worth noting that in order to describe DWARF reasoning accuracy in the whole training process, Entity State Similarity (ESS, be similar to CSS we mentioned above) is used as an evaluation index.
Attention Mechanism
We first focus on ablation experiments on the attention mechanism. As shown in Figure 10, the absence of the attention mechanism has little effect on WN18RR (sparse KG), while the effect of the attention mechanism increases with increasing density. In particular, on FB15K-237 (dense KG), the effect of the attention mechanism on the average ESS reaches $1/3$. It is consistent with the purpose of introducing the attention mechanism, that is, enhancing the reasoning ability of DWARF in the case of multi-neighbor and multi-relation. The reason lies in DWARF are forced to prioritize relations and neighbors that are highly correlated with the query relations, which plays an important role in reasoning on dense KG with complex relations, but because sparse KG neighbor relations are not sophisticated, the attention mechanism does not have a substantial effect.
Figure 10: Entity state similarities comparing the performance of FULORA, FULORA-GE, FULORA-ATTN and FULORA-DPF. In addition, bar charts and distribution charts depict the learning speed and accuracy. Learning curves are averaged over 3 seeds.
Dynamic Path Feedback
We have demonstrated in previous experiments that dynamic path feedback can improve GIANT’s reasoning ability, next we concentrate on the impact to DWARF’s learning speed of dynamic path feedback. To illustrate the speed and stability of the training, we utilize kdeplot in Figure 10. The distribution maps on the three datasets have a common feature: compared with FULORA, the learning speed and learning accuracy decrease in the absence of dynamic path feedback (corresponding to the widening and downward movement of the distribution maps). In FB15K-237 (dense KG), the lack of dynamic path feedback has a great impact on the learning speed and accuracy, because the error of cluster mapping escalates with the complexity of neighbor relations when the number of clusters maintains. At this time, a large deviation between cluster-level KG and entity-level KG requires GIANT to make a correct evaluation of the path timely to provide exact guidance for DWARF.
Efficient Guidance-Exploration
The core of FULORA is the efficient guidance-exploration method, which contributes to make the guidance-exploration trade-off. In the previous HRL-based reasoning method like CURL, there is a high degree of coupling between the two agents. We treat ESS and CSS as two time series because they change with training progress. Next, we use Granger Causality test to examine the degree of coupling between DWARF and GIANT.
Granger Causality test is used to study the causal relationship between two sets of data, that is, to test whether one set of time series causes changes in another set of time series (Diks and Panchenko 2006). It is worth noting that Granger Causality test requires stationary time series, otherwise false regression issues may occur, so it is necessary to detect the stationarity of time series through ADF test. If the pair-to-pair time series is non-stationary and satisfies the homogeneity of order, the Granger Causality test can be carried out only after the cointegration test between pair-to-pair sequences exists (Dougherty 2011). Our time series analysis for ESS and CSS are detailed in Appendix C.
The results of Granger Causality test for FULORA and FULORA-GE are shown in Table 6. Efficient guidance-exploration method reduces the coupling between GIANT and DWARF, thereby alleviating the affect of distribution deviation caused by cluster mapping. In particular, on the WN18RR and FB15K-237, FULORA performs better than the version without efficient guidance-exploration method due to the greatly reduced coupling between DWARF and GIANT. Certain tasks in WN18RR (sparse KG) require the agent to have excellent long-distance reasoning ability. However, due to the distribution deviation, even if the GIANT reaches the correct target cluster, the guidance provided by it may not be suitable for DWARF (for example, the path length is too long to reach the correct target entity). The high coupling causes GIANT also seriously affected by DWARF, resulting in poor model performance.
Relation NELL-995 WN18RR FB15K-237
DWARF $\rightarrow$ GIANT 21.5 9.6 18.9
DWARF(-GE) $\rightarrow$ GIANT(-GE) 25.3 33.9 60.4
GIANT $\rightarrow$ DWARF 21.7 40.3 36.3
GIANT(-GE) $\rightarrow$ DWARF(-GE) 26.3 41.1 37.3
Table 6: F statistics of Granger Causality test for FULORA and FULORA-GE. DWARF (-GE) and GIANT (-GE) are from FULORA-GE. In the Granger Causality test, a large F statistic indicates a strong causal relation. We set the lagged step as 2.
Appendix C C. Time Series Analysis to Efficient Guidance-Exploration
Before Granger Causality test, we firstly use Augmented Dickey-Fuller (ADF) test to judge the stationarity of the series. The results as shown in Table 7. Next, we make D-order difference for the non-stationary time series, and then test the stationarity of the difference series. As illustrated in Figure 11, the time series is stationary after 1-order difference. Therefore, we demonstrate that these series are integrated of order one. To avoid the pseudo-regression phenomenon, we also need to do cointegration analysis of these series (Dougherty 2011). Here, we employ Johansen test based on maximum likelihood estimation for cointegration analysis.
Figure 11: Entity state similarity curves, cluster state similarity curves and their 1-order difference on WN18RR and FB15K-237 from FULORA and FULORA-GE.
In the Johansen cointegration test, the 2-cointegration term critical values of {10%, 5%, 1%} are {2.705, 3.841, 6.635} respectively. The critical value 3.841 corresponding to 5% was selected as the boundary for testing. Table 8 shows the results of Johansen cointegration test for FULORA and FULORA-GE. It shows that the seven non-stationary series avoid the pseudo-regression problem, thus the result of Granger causality test is reasonable.
Variable NELL-995 WN18RR FB15K-237
$t$ $P$ $t$ $P$ $t$ $P$
DWARF -3.25 0.04 -7.40 0.00 -1.67 0.45
GIANT -5.22 0.00 -1.46 0.46 0.08 0.97
DWARF(-GE) -3.11 0.04 -1.65 0.46 0.08 0.99
GIANT(-GE) -4.14 0.00 -1.98 0.30 -0.21 0.94
Table 7: t statistics and P value of ADF test for FULORA and FULORA-GE. DWARF (-GE) and GIANT (-GE) are from FULORA-GE. We set the bound of P value to 0.05, and P-values above the bound are bolded to indicate that they are not stationary series.
Relation WN18RR FB15K-237
CR Trace CR Trace
DWARF $\leftrightarrow$ GIANT 0.019 38.436 0.002 4.736
DWARF(-GE) $\leftrightarrow$ GIANT(-GE) 0.037 74.582 0.002 4.133
Table 8: Results of Johansen cointegration test. CR is the abbreviation of Characteristic Root. The trace satisfying the critical value is bolded, indicating that there are two cointegration relations.
Appendix D D. Proof of Dynamic Path Feedback
A major concern with dynamic path feedback is whether the optimal policy GIANT learns is consistent with the default rewards. Recall Eq.(12) where the next state $s^{c}_{t+1}$ is used, which means that it implies the action $a^{c}_{t}$. The next state $s^{c}_{t+1}$ is the output of the current state $s^{c}_{t}$ and the current action $a^{c}_{t}$ to the transition function. In fact, $J(\theta_{\pi^{c}})$ is a function of state $s^{c}_{t}$ and action $a^{c}_{t}$. The same goes for reward settings in the default environment. Therefore, we use the Q-function to simplify our notation,
$Q(s^{c}_{t},a^{c}_{t})=\sum_{\tau=t}^{T-1}[r_{c}(s_{\tau}^{c})].$ (13)
The optimal Q-function $Q^{*}(s^{c}_{t},a^{c}_{t})$ is subject to Bellman optimal equation (Bellman 1958):
$Q^{*}(s_{t}^{c},a_{t}^{c})$ (14)
$=\mathbb{E}_{s_{t+1}^{c}\sim P(\cdot|s_{t}^{c},a_{t}^{c})}\left[r_{c}(s_{t}^{c})+\underset{a_{t+1}^{c}\in\mathcal{A}}{\text{max}}Q^{*}(s^{c}_{t+1},a^{c}_{t+1})\right],$
then we make a simple transformation of the above formula to get
$Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\text{target}})$ (15)
$=\mathbb{E}_{s_{t+1}^{c}}\left[r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})\right.$
$\left.+\underset{a_{t+1}^{c}\in\mathcal{A}}{\text{max}}(Q^{*}(s^{c}_{t+1},a^{c}_{t+1})-\alpha D(s^{c}_{t+1},s^{c}_{\text{target}}))\right].$
The $D(s^{c}_{t},s^{c}_{\text{target}})$ we define is only related to the state, so Eq.(15) is equivalent to Eq.(14). Notice that $r_{c}(s_{t}^{c})-\alpha\Delta(s_{t}^{c},s_{t+1}^{c})$ is the new reward function $\hat{r}_{c}(s^{c}_{t})$ we designed for GIANT. Hence, Eq.(15) can be viewed as
$\hat{Q}^{*}(s_{t}^{c},a_{t}^{c})$ (16)
$=\mathbb{E}_{s_{t+1}^{c}}\left[\hat{r}_{c}(s^{c}_{t})+\underset{a_{t+1}^{c}\in\mathcal{A}}{\text{max}}\hat{Q}^{*}(s^{c}_{t+1},a^{c}_{t+1})\right],$
where $\hat{Q}^{*}(s^{c}_{t},a^{c}_{t})=Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\text{target}})$. Eq.(16) is the Bellman optimal equation in dynamic path feedback. It constructs a new Markov Decision Process (MDP), which we use $M^{\prime}$ to denote for distinction and $M$ for the original MDP. Thus, the optimal policy in $M^{\prime}$ is
$\pi^{*}_{M^{\prime}}(s_{t}^{c})$ $=\text{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\text{max}}\hat{Q}^{*}(s_{t}^{c},a_{t}^{c})$ (17)
$=\text{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\text{max}}Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\text{target}})$
$=\text{arg}\underset{a_{t}^{c}\in\mathcal{A}}{\text{max}}Q^{*}(s_{t}^{c},a_{t}^{c})=\pi^{*}_{M}(s_{t}^{c}),$
which implies that GIANT learns optimal policy in dynamic path feedback is consistent with optimalcpolicy in the default rewards.